OpenStudy (anonymous):

Algebraic equation:-1 equals 5+x over 6. the 6 is over the whole thing. thx so much for ur help

OpenStudy (anonymous):

\[-1=\frac{5+x}{6}\] \[-1\times6=5+x\] \[-6=5+x\] \[x=-6-5\] \[x=-11\]

OpenStudy (anonymous):

thanks sooooooo much that was really helpful